An increase in the demand for green tea raises the price of green tea from $16 a pound to $20 a pound. As a result, quantity supplied increases by 30 percent. Using the midpoint formula, what is the value of the price elasticity of supply?
Answer #1

Given p1=$16;p2=$20

%change in quantity supply=30%

By using the mid point formula

Price elasticity of supply=%change in quantity supply/%change in price

%change in price=(p2-p1÷p2+p1/2)×100

=($20-$16÷$20+$16/2)×100

=(4/18)×100=22.22%

Price elasticity of supply=30%/22.22%=1.35